# Pre-supernova O-C shell mergers could produce more 44Ti than the explosion

### v2

@@ Abstract
Context. The formation of $^{44}\mathrm{Ti}$ in massive stars is thought to occur during explosive nucleosynthesis, however recent studies have shown it can be produced during O-C shell mergers prior to core collapse.
Aims. We investigate how mixing according to 3D macro physics derived from hydrodynamic simulations impacts the pre-supernova O-C shell merger nucleosynthesis and if it can dominate the explosive supernova production of $^{44}\mathrm{Ti}$ and other radioactive isotopes.
Methods. We compare a range of observations and models of explosive $^{44}\mathrm{Ti}$ yields to pre-explosive multi-zone mixing-burning nucleosynthesis simulations of an O-C shell merger in a $15\penalty 10000\ \mathrm{M_{\odot}}$ stellar model with mixing conditions corresponding to different 3D hydro mixing scenarios.
Results. Radioactive species produced in the $\mathrm{O}$ shell have a spread in their pre-explosive yields predictions across different 3D mixing scenarios of $2.14\penalty 10000\ \mathrm{dex}$ on average. $^{44}\mathrm{Ti}$ has the largest spread of $4.78\penalty 10000\ \mathrm{dex}$. The pre-explosive production of $^{44}\mathrm{Ti}$ can be larger than the production of all massive star models in the NuGrid data set where ${{\mathrm{}}}{{\mathrm{{}^{44}}}}{\mathrm{Ti}}$ is dominated by the explosive nucleosynthesis contribution, as well all other massive star and supernova models.
Conclusions. We conclude that quantitative predictions of $^{44}\mathrm{Ti}$ and other radioactive species more broadly require an understanding of the 3D hydrodynamic mixing conditions present during the O-C shell merger.
Key Words.:
convection – nuclear reactions, nucleosynthesis, abundances – stars: massive – stars: interiors – supernovae: general
@@
1 Introduction
@@
The radioactive isotope $\mathrm{{}^{44}Ti}$ is a key tracer of explosive nucleosynthesis in core collapse supernovae (CCSN). It is primarily synthesized during $\alpha$-rich freeze-out from nuclear statistical equilibrium in the innermost ejecta associated with explosive $\mathrm{Si}$ burning by the reactions $\mathrm{{}^{40}Ca}$$(\alpha,\gamma)$ and $\mathrm{{}^{43}Sc}$$(p,\gamma)$ (Woosley & Weaver, 1995; The et al., 2006; Magkotsios et al., 2010; Sieverding et al., 2023). It is observed by the emission lines $67.87$, $78.32$, and $1157\;\mathrm{\mathrm{k}\mathrm{eV}}$ from the decay chain $\mathrm{{}^{44}Ti}\rightarrow\mathrm{{}^{44}Sc}\rightarrow\mathrm{{}^{44}Ca}$ in supernova remnants (Iyudin et al., 1994; Grebenev et al., 2012; Boggs et al., 2015; Grefenstette et al., 2014, 2017).
@@
Recently it has been shown that $\mathrm{{}^{44}Ti}$ can be significantly produced prior to the explosion during an $\mathrm{O}$-$\mathrm{C}$ shell merger in massive stars (Roberti et al., 2025). Mergers occur a few hours prior to the CCSN of a red supergiant, therefore $\mathrm{{}^{44}Ti}$ ($t_{1/2}=59.1\;\mathrm{\mathrm{yr}}$) synthesized in this event can contribute to the observed signal seen in supernova remnants. This may be the case for Cassiopeia A, as observations of its $\mathrm{Si}$, $\mathrm{Ne}$, $\mathrm{Cl}$, and $\mathrm{P}$ have recently been interpreted as signatures of an $\mathrm{O}$-$\mathrm{C}$ shell merger prior to collapse (Sato et al., 2025; Audard et al., 2025).
@@
Cassiopeia A (Iyudin et al., 1994; Grefenstette et al., 2014, 2017) and SN 1987A (Grebenev et al., 2012; Boggs et al., 2015) both have asymmetric and clumpy distributions of $\mathrm{{}^{44}Ti}$. This may be a signature of the pre-explosive origin of $\mathrm{{}^{44}Ti}$ in an $\mathrm{O}$-$\mathrm{{}^{12}C}$shell merger, as 3D hydrodynamic simulations show mergers have asymmetric and large-scale non-radial flows whose features could be preserved in the remnant (Andrassy et al., 2020; Yadav et al., 2020; Rizzuti et al., 2024; Sato et al., 2025). Nucleosynthesis during mergers in these 3D macro physical flows diverges significantly from 1D stellar evolution predictions in the $\mathrm{O}$ shell (Yadav et al., 2020; Rizzuti et al., 2024; Issa et al., 2025a, b).
@@
In this study, we explore how pre-explosive nucleosynthesis of $\mathrm{{}^{44}Ti}$ and other radioactive isotopes depend on the mixing conditions during an $\mathrm{O}$-$\mathrm{C}$ shell merger as motivated by 3D hydrodynamic simulations. This Letter is organized as follows: §2 provides an overview of the post-processed O-C shell conditions, §3 demonstrates how $\mathrm{{}^{44}Ti}$ is produced non-explosively and compares the production in the $\mathrm{O}$-$\mathrm{C}$ shell merger to explosive yields, and finally in §4 we discuss the implications of these results for observations.
2 Methods
@@
The NuGrid data set includes massive star models with initial masses $12{-}25\;\mathrm{\mathrm{M}_{\odot}}$ and initial metallicities $Z=10^{-4}$ to $0.02$ (Ritter et al., 2018). These models were calculated in 1D using MESA (Paxton et al., 2010) without rotation and no convective overshooting after the end of core helium burning (Pignatari et al., 2016; Ritter et al., 2018). Four models have an $\mathrm{O}$-$\mathrm{C}$ shell merger shortly before core collapse: $15\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.02$, $12\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.01$, $15\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.01$, and $20\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.02$.
@@
The mixing scenarios inspired by 3D hydrodynamic simulations (Figures 4 and 5, Issa et al., 2025a) include a downturn in the mixing efficiency profile at the bottom of the $\mathrm{O}$ shell (Meakin & Arnett, 2006; Jones et al., 2017), increased convective velocities compared to mixing length theory (MLT) of factors 33, $10$, and $50$ (a range seen in the simulations of Jones et al., 2017; Andrassy et al., 2020; Rizzuti et al., 2024), and quenching effects due to possibly strong energy feedback and partial merging (Andrassy et al., 2020). They also have a range of $\mathrm{C}$-shell entrainment rates into the $\mathrm{O}$ shell from no entrainment to a full merger of the convective $\mathrm{O}$ and $\mathrm{C}$ shells, as the simulations of Andrassy et al. (2020) show that entrainment depends on the luminosity of $\mathrm{{}^{12}C}+\mathrm{{}^{12}C}$ burning.
@@
$\mathrm{OP}_{\mathrm{R}15}=X_{\mathrm{model}}/X_{\mathrm{Pre-Explosive\penalty 10000\ 15\;\mathrm{M_{\odot}}\penalty 10000\ Z=0.02}}$
where $X_{\mathrm{model}}$ is the ejected yield for a species from a given model and $X_{\mathrm{Pre-Explosive\penalty 10000\ 15\;\mathrm{M_{\odot}}\penalty 10000\ Z=0.02}}$ is the undecayed pre-explosive yield from the $15\;\mathrm{\mathrm{M}_{\odot}}$, $Z=0.02$ model from Ritter et al. (2018). The simulations from Issa et al. (2025a, b) simulate the $\mathrm{O}$ shell alone for $110\;\mathrm{\mathrm{s}}$, and so the yields for these models are estimated according to the method described by §2.3 of Issa et al. (2025b). Further simulation details can be found in Issa et al. (2025a, b).
3 Results
3.1 Nuclear production of $\mathrm{{}^{44}Ti}$ during O-C shell mergers
@@
The mixing cases with a $50\times D_{\mathrm{3D{-}insp.}}$ profile have some exceptions to the $\mathrm{{}^{38}Ar}$ and $\mathrm{{}^{42}Ca}$ contributions to $\mathrm{{}^{44}Ti}$. These cases feature significant contributions of $\mathrm{{}^{39}K}$$(\alpha,p)$$\mathrm{{}^{42}Ca}$ and $\mathrm{{}^{43}Sc}$$(\gamma,p)$$\mathrm{{}^{42}Ca}$ along with $\mathrm{{}^{38}Ar}$$(\alpha,\gamma)$$\mathrm{{}^{42}Ca}$. The contribution of $\mathrm{{}^{43}Sc}$ to $\mathrm{{}^{42}Ca}$ may seem counter-intuitive as Fig. 2 shows $\mathrm{{}^{42}Ca}$$(p,\gamma)$$\mathrm{{}^{43}Sc}$ as a destruction channel for $\mathrm{{}^{42}Ca}$. However, it is because the $\mathrm{O}$ shell is a convective-reactive enviornment where species can advect on a similar timescale to their reactions. Fig. 2 shows a build-up of $\mathrm{{}^{43}Sc}$ which is mixed deeper in the $\mathrm{O}$ and undergoes $\mathrm{{}^{43}Sc}$$(\gamma,p)$$\mathrm{{}^{42}Ca}$ at those hotter temperatures. For $\mathrm{{}^{38}Ar}$, in all other cases it is produced by $\mathrm{{}^{30}Si}$$(\alpha,\gamma)$$\mathrm{{}^{34}S}$$(\alpha,\gamma)$$\mathrm{{}^{38}Ar}$, but in the $50\times D_{\mathrm{3D{-}insp.}}$ mixing cases it is also produced by $\mathrm{{}^{35}Cl}$$(\gamma,p)$$\mathrm{{}^{34}S}$$(\alpha,\gamma)$$\mathrm{{}^{38}Ar}$ and $\mathrm{{}^{39}K}$$(\gamma,p)$$\mathrm{{}^{38}Ar}$. How $\mathrm{{}^{30}Si}$ and the light odd-Z isotopes $\mathrm{{}^{35}Cl}$ and $\mathrm{{}^{39}K}$ are produced in the $\mathrm{O}$-$\mathrm{C}$ shell merger is discussed in Issa et al. (2025b). The importance of the $(\alpha,\gamma)$ channels of $\mathrm{{}^{34}S}$ and $\mathrm{{}^{38}Ar}$ for $\mathrm{{}^{44}Ti}$ production are interesting as their $(p,\gamma)$ channels may dominate energy production in the $\mathrm{O}$-$\mathrm{C}$ shell (Roberti & Pignatari, 2025).
@@
3.2 Mixing and explosive yields of radioactive species
@@
Figure 4: $\mathrm{OP}_{\mathrm{R15}}$ for (1) inferred yields of $\mathrm{{}^{44}Ti}$ from supernova remnants observations (triangles with error bars) (2) explosive $\mathrm{{}^{44}Ti}$ yields of non-NuGrid models (blue circles) (3) explosive yields of all considered radioactive species from NuGrid models (small black stars) (4) explosive yields of the $15\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.02$ NuGrid model (large black stars) (5) the maximum and minimum predicted pre-explosive yields for all scenarios (purple bars). Appendix A provides the full list of citations.
@@
The range of predicted pre-explosive yields in our mixing scenarios for $\mathrm{{}^{22}Na}$, $\mathrm{{}^{44}Ti}$, and $\mathrm{{}^{51}Cr}$ are comparable or larger than the explosive yields of all massive star models in the NuGrid data set. Our pre-explosive yields of $\mathrm{{}^{59}Fe}$, $\mathrm{{}^{60}Fe}$, and $\mathrm{{}^{65}Zn}$ are also comparable or larger than many of the explosive yields in the NuGrid data set. Table 2 in Appendix A shows that explosive nucleosynthesis dominates $\mathrm{{}^{44}Ti}$ yields in the NuGrid data set except for the $15\;\mathrm{\mathrm{M}_{\odot}}$, $Z=0.02$ model and the $20\;\mathrm{\mathrm{M}_{\odot}}$, $Z=0.01$ model as the only two models where pre-explosive yields of $\mathrm{{}^{44}Ti}$ are comparable to the explosion. Notably the other two models with $\mathrm{O}$-$\mathrm{C}$ shell mergers do not show significant pre-explosive production of $\mathrm{{}^{44}Ti}$, but could do so if 3D mixing effects were included as in this work.
We have also compared our results to a number of 1D, 2D, and 3D models of massive star yields of $\mathrm{{}^{44}Ti}$ from the literature, some of which include rotation. Fig. 4 shows that the range of predicted pre-explosive yields for $\mathrm{{}^{44}Ti}$ can dominate regardless of the different modelling assumptions and dimensionality of the explosion. Additionally, our predicted range does not exceed the inferred yields of $\mathrm{{}^{44}Ti}$ from observed supernova remnants, although the largest upper limit comes from Vela Junior which has large uncertainties in its distance and age (Weinberger et al., 2020). Even so, our results are not inconsistent with observational evidence. For instance, according to Weinberger et al. (2020) Cassiopeia A ejected $(2.6\pm 0.6)\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}}$ of $\mathrm{{}^{44}Ti}$ which is between our $3\times$ and $10\times D_{\mathrm{3D{-}insp.}}$ cases with the fastest entrainment rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. If the progenitor of Cassiopeia A underwent an $\mathrm{O}$-$\mathrm{C}$ shell merger prior to explosion (Sato et al., 2025; Audard et al., 2025), the mixing profile of its $\mathrm{O}$ shell having a convective downturn and mixing speeds boosted by 3103{-}10 compared to MLT would explain the $\mathrm{{}^{44}Ti}$ yield without much synthesis in the explosion. These results demonstrate the critical importance of understanding the mixing conditions during the $\mathrm{O}$-$\mathrm{C}$ shell merger for predicting the yields of radioactive isotopes for pre-supernova models, especially $\mathrm{{}^{44}Ti}$.
4 Conclusions

### v4

@@ Abstract
The formation of $^{44}\mathrm{Ti}$ in massive stars is thought to occur during explosive nucleosynthesis, however recent studies have shown it can be produced during O-C shell mergers prior to core collapse. We investigate how mixing according to 3D macro physics derived from hydrodynamic simulations impacts the pre-supernova O-C shell merger nucleosynthesis and if it can dominate the explosive supernova production of $^{44}\mathrm{Ti}$ and other radioactive isotopes. We compare a range of observations and models of explosive $^{44}\mathrm{Ti}$ yields to pre-explosive multi-zone mixing-burning nucleosynthesis simulations of an O-C shell merger in a $15~\mathrm{M_{\odot}}$ stellar model with mixing conditions corresponding to different 3D hydro mixing scenarios. Radioactive species produced in the $\mathrm{O}$ shell have a spread in their pre-explosive yields predictions across different 3D mixing scenarios of $2.14~\mathrm{dex}$ on average. $^{44}\mathrm{Ti}$ has the largest spread of $4.78~\mathrm{dex}$. The pre-explosive production of $^{44}\mathrm{Ti}$ can be larger than the production of all massive star models in the NuGrid data set where $^{44}\mathrm{Ti}$ is dominated by the explosive nucleosynthesis contribution, as well all other massive star and supernova models. We conclude that quantitative predictions of $^{44}\mathrm{Ti}$ and other radioactive species more broadly require an understanding of the 3D hydrodynamic mixing conditions present during the O-C shell merger.
@@
journal: ApJL
show]joshuaissa@uvic.ca
I Introduction
@@
The radioactive isotope $\mathrm{{}^{44}Ti}$ is a key tracer of explosive nucleosynthesis in core collapse supernovae (CCSN). It is primarily synthesized during $\alpha$-rich freeze-out from nuclear statistical equilibrium in the innermost ejecta associated with explosive $\mathrm{Si}$ burning by the reactions $\mathrm{{}^{40}Ca}$$(\alpha,\gamma)$ and $\mathrm{{}^{43}Sc}$$(p,\gamma)$ (Woosley and Weaver, 1995; The et al., 2006; Magkotsios et al., 2010; Sieverding et al., 2023). It is observed by the emission lines $67.87$, $78.32$, and $1157\;\mathrm{\mathrm{k}\mathrm{eV}}$ from the decay chain $\mathrm{{}^{44}Ti}\rightarrow\mathrm{{}^{44}Sc}\rightarrow\mathrm{{}^{44}Ca}$ in supernova remnants (Iyudin et al., 1994; Grebenev et al., 2012; Boggs et al., 2015; Grefenstette et al., 2014, 2017).
@@
Recently it has been shown that $\mathrm{{}^{44}Ti}$ can be significantly produced prior to the explosion during an $\mathrm{O}$-$\mathrm{C}$ shell merger in massive stars (Tur et al., 2010; Chieffi and Limongi, 2017; Roberti et al., 2025). Mergers occur a few hours prior to the CCSN of a red supergiant, therefore $\mathrm{{}^{44}Ti}$ ($t_{1/2}=59.1\;\mathrm{\mathrm{yr}}$) synthesized in this event can contribute to the observed signal seen in supernova remnants. This may be the case for Cassiopeia A, as observations of its $\mathrm{Si}$, $\mathrm{Ne}$, $\mathrm{Cl}$, and $\mathrm{P}$ have recently been interpreted as signatures of an $\mathrm{O}$-$\mathrm{C}$ shell merger prior to collapse (Sato et al., 2025; Audard et al., 2025).
@@
Cassiopeia A (Iyudin et al., 1994; Grefenstette et al., 2014, 2017) and SN 1987A (Grebenev et al., 2012; Boggs et al., 2015) both have asymmetric and clumpy distributions of $\mathrm{{}^{44}Ti}$. This may be a signature of the pre-explosive origin of $\mathrm{{}^{44}Ti}$ in an $\mathrm{O}$-$\mathrm{C}$ shell merger, as 3D hydrodynamic simulations show mergers have asymmetric and large-scale non-radial flows whose features could be preserved in the remnant (Andrassy et al., 2020; Yadav et al., 2020; Rizzuti et al., 2024; Sato et al., 2025). Nucleosynthesis during mergers in these 3D macro physical flows diverges significantly from 1D stellar evolution predictions in the $\mathrm{O}$ shell (Yadav et al., 2020; Rizzuti et al., 2024; Issa et al., 2025a, b).
@@
In this study, we explore how pre-explosive nucleosynthesis of $\mathrm{{}^{44}Ti}$ and other radioactive isotopes depend on the mixing conditions during an $\mathrm{O}$-$\mathrm{C}$ shell merger as motivated by 3D hydrodynamic simulations. This Letter is organized as follows: §II provides an overview of the post-processed O-C shell conditions, §III demonstrates how $\mathrm{{}^{44}Ti}$ is produced non-explosively and compares the production in the $\mathrm{O}$-$\mathrm{C}$ shell merger to explosive yields, and finally in §IV we discuss the implications of these results for observations.
II Methods
@@
The NuGrid data set includes massive star models with initial masses $12{-}25\;\mathrm{\mathrm{M}_{\odot}}$ and initial metallicities $Z=10^{-4}$ to $0.02$ (Ritter et al., 2018). These models were calculated in 1D using MESA (Paxton et al., 2010) without rotation and no convective overshooting after the end of core helium burning (Pignatari et al., 2016; Ritter et al., 2018). Five models have an $\mathrm{O}$-$\mathrm{C}$ shell merger shortly before core collapse: $15\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.02$, $12\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.01$, $15\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.01$, $20\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.01$, and $12\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.001$.
@@
The mixing scenarios inspired by 3D hydrodynamic simulations (Figures 4 and 5, Issa et al., 2025a) include a downturn in the mixing efficiency profile at the bottom of the $\mathrm{O}$ shell ($D_{\mathrm{3D{-}insp.}}$ in this work, inspired by Meakin and Arnett, 2006; Jones et al., 2017), increased convective velocities compared to mixing length theory (MLT) of factors 33, $10$, and $50$ (a range seen in the simulations of Jones et al., 2017; Andrassy et al., 2020; Rizzuti et al., 2024), and quenching effects due to possibly strong energy feedback and partial merging (Andrassy et al., 2020). They also have a range of $\mathrm{C}$-shell entrainment rates into the $\mathrm{O}$ shell from no entrainment to a full merger of the convective $\mathrm{O}$ and $\mathrm{C}$ shells, as the simulations of Andrassy et al. (2020) show that entrainment depends on the luminosity of $\mathrm{{}^{12}C}+\mathrm{{}^{12}C}$ burning.
@@
$\mathrm{OP}_{\mathrm{R}15}=X_{\mathrm{model}}/X_{\mathrm{Pre-Explosive~15\;\mathrm{M_{\odot}}~Z=0.02}}$
where $X_{\mathrm{model}}$ is the ejected yield for a species from a given model and $X_{\mathrm{Pre-Explosive~15\;\mathrm{M_{\odot}}~Z=0.02}}$ is the undecayed pre-explosive yield from the $15\;\mathrm{\mathrm{M}_{\odot}}$, $Z=0.02$ model from Ritter et al. (2018). The simulations from Issa et al. (2025a, b) simulate the $\mathrm{O}$ shell alone for $110\;\mathrm{\mathrm{s}}$, and so the yields for these models are estimated according to the method described by §2.3 of Issa et al. (2025b). Further simulation details can be found in Issa et al. (2025a, b).
III Results
III.1 Nuclear production of $\mathrm{{}^{44}Ti}$ during O-C shell mergers
@@
The mixing cases with a $50\times D_{\mathrm{3D{-}insp.}}$ profile have some exceptions to the $\mathrm{{}^{38}Ar}$ and $\mathrm{{}^{42}Ca}$ contributions to $\mathrm{{}^{44}Ti}$. These cases feature significant contributions of $\mathrm{{}^{39}K}$$(\alpha,p)$$\mathrm{{}^{42}Ca}$ and $\mathrm{{}^{43}Sc}$$(\gamma,p)$$\mathrm{{}^{42}Ca}$ along with $\mathrm{{}^{38}Ar}$$(\alpha,\gamma)$$\mathrm{{}^{42}Ca}$. The contribution of $\mathrm{{}^{43}Sc}$ to $\mathrm{{}^{42}Ca}$ may seem counter-intuitive as Fig. 2 shows $\mathrm{{}^{42}Ca}$$(p,\gamma)$$\mathrm{{}^{43}Sc}$ as a destruction channel for $\mathrm{{}^{42}Ca}$. However, it is because the $\mathrm{O}$ shell is a convective-reactive enviornment where species can advect on a similar timescale to their reactions. Fig. 2 shows a build-up of $\mathrm{{}^{43}Sc}$ which is mixed deeper in the $\mathrm{O}$ and undergoes $\mathrm{{}^{43}Sc}$$(\gamma,p)$$\mathrm{{}^{42}Ca}$ at those hotter temperatures. For $\mathrm{{}^{38}Ar}$, in all other cases it is produced by $\mathrm{{}^{30}Si}$$(\alpha,\gamma)$$\mathrm{{}^{34}S}$$(\alpha,\gamma)$$\mathrm{{}^{38}Ar}$, but in the $50\times D_{\mathrm{3D{-}insp.}}$ mixing cases it is also produced by $\mathrm{{}^{35}Cl}$$(\gamma,p)$$\mathrm{{}^{34}S}$$(\alpha,\gamma)$$\mathrm{{}^{38}Ar}$ and $\mathrm{{}^{39}K}$$(\gamma,p)$$\mathrm{{}^{38}Ar}$. How $\mathrm{{}^{30}Si}$ and the light odd-Z isotopes $\mathrm{{}^{35}Cl}$ and $\mathrm{{}^{39}K}$ are produced in the $\mathrm{O}$-$\mathrm{C}$ shell merger is discussed in Issa et al. (2025b). The importance of the $(\alpha,\gamma)$ channels of $\mathrm{{}^{34}S}$ and $\mathrm{{}^{38}Ar}$ for $\mathrm{{}^{44}Ti}$ production are interesting as their $(p,\gamma)$ channels may dominate energy production in the $\mathrm{O}$-$\mathrm{C}$ shell (Roberti and Pignatari, 2025).
@@
III.2 Mixing and explosive yields of radioactive species
@@
Figure 4: $\mathrm{OP}_{\mathrm{R15}}$ for (1) inferred yields of $\mathrm{{}^{44}Ti}$ from supernova remnants observations (triangles with error bars) (2) explosive $\mathrm{{}^{44}Ti}$ yields of non-NuGrid models (blue circles) (3) explosive yields of all considered radioactive species from NuGrid models (small black stars) (4) explosive yields of the $15\;\mathrm{\mathrm{M}_{\odot}}$ $Z=0.02$ NuGrid model (large black stars) (5) the maximum and minimum predicted pre-explosive yields for all scenarios (purple bars). Appendix B provides the full list of citations.
@@
The range of predicted pre-explosive yields in our mixing scenarios for $\mathrm{{}^{22}Na}$, $\mathrm{{}^{44}Ti}$, and $\mathrm{{}^{51}Cr}$ are comparable or larger than the explosive yields of all massive star models in the NuGrid data set. Our pre-explosive yields of $\mathrm{{}^{59}Fe}$, $\mathrm{{}^{60}Fe}$, and $\mathrm{{}^{65}Zn}$ are also comparable or larger than many of the explosive yields in the NuGrid data set. Table 2 in Appendix B shows that explosive nucleosynthesis dominates $\mathrm{{}^{44}Ti}$ yields in the NuGrid data set except for the $15\;\mathrm{\mathrm{M}_{\odot}}$, $Z=0.02$ model and the $20\;\mathrm{\mathrm{M}_{\odot}}$, $Z=0.01$ model as the only two models where pre-explosive yields of $\mathrm{{}^{44}Ti}$ are comparable to the explosion. Notably the other three models with $\mathrm{O}$-$\mathrm{C}$ shell mergers do not show significant pre-explosive production of $\mathrm{{}^{44}Ti}$, but could do so if 3D mixing effects were included as in this work.
We have also compared our results to a number of 1D, 2D, and 3D models of massive star yields of $\mathrm{{}^{44}Ti}$ from the literature, some of which include rotation. Fig. 4 shows that the range of predicted pre-explosive yields for $\mathrm{{}^{44}Ti}$ can dominate regardless of the different modelling assumptions and dimensionality of the explosion. Additionally, our predicted range does not exceed the inferred yields of $\mathrm{{}^{44}Ti}$ from observed supernova remnants, although the largest upper limit comes from Vela Junior which has large uncertainties in its distance and age (Weinberger et al., 2020). Comparing our results to Weinberger et al. (2020) Cassiopeia A ejected $(2.6\pm 0.6)\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}}$ of $\mathrm{{}^{44}Ti}$ which is between our $3\times$ and $10\times D_{\mathrm{3D{-}insp.}}$ cases with the fastest entrainment rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. If the progenitor of Cassiopeia A underwent an $\mathrm{O}$-$\mathrm{C}$ shell merger prior to explosion (Sato et al., 2025; Audard et al., 2025), the mixing profile of its $\mathrm{O}$ shell having a convective downturn and mixing speeds boosted by 3103{-}10 compared to MLT would explain the $\mathrm{{}^{44}Ti}$ yield without much synthesis in the explosion. These results demonstrate the critical importance of understanding the mixing conditions during the $\mathrm{O}$-$\mathrm{C}$ shell merger for predicting the yields of radioactive isotopes for pre-supernova models, especially $\mathrm{{}^{44}Ti}$.
IV Conclusions